OpenStudy (anonymous):

1. (g + 4) – 3g = 1 + g (1 point

OpenStudy (anonymous):

g+4-3g = 1 + g -3g = -3 g =1

OpenStudy (anonymous):

how do i solve this problum

OpenStudy (lalaly):

g+4-3g=1+g -2g+4=1+g -3g=-3 g=1